# Curriculum effects and compositionality emerge with in-context learning in neural networks

### v3

@@ 2.2 Curriculum effects in a compositional task
Figure 3: Compositional task and results. The task is derived from a human study [19]. Networks were presented with the locations corresponding to particular cues (colored animals) and had to predict the locations of unseen cues. Cues were comprised of sequences of tokens (e.g., ‘blue alligator’). See Appendix for details. (a) Curriculum conditions. In both the blocked and interleaved conditions, the 9 study examples always included one full row and one full column. In the blocked condition, the row was presented in full before the column, or vice-versa. In the interleaved condition, these 9 examples were randomly shuffled. (b) In the rule-like condition, locations varied systematically with the color and animal features (e.g., color determined x-coordinate and animal determined y-coordinate). (c) In our novel rotated condition, the original axes were rotated by 45 degrees, so that any change in either color or animal resulted in a change to both x- and y-coordinates. (d-f) Performance on the compositional task with in-weight learning (IWL) only, where again randomly initialized networks were trained from scratch. (d) The few-shot evaluation tested networks’ ability to solve the task in context based on the 9 study examples given in the input. Again, without prior metalearning neural networks were incapable of solving the task in this way, regardless of condition. (e) Without prior metalearning experience, the network was still able to learn via IWL, performing well on both the rule-like and the rotated tasks after finetuning. IWL again exhibited an interleaving advantage due to catastrophic forgetting (see f). (f) Accuracy and loss results over the course of IWL training in each of the four conditions. Accuracy is again split by group, in this case corresponding to whether the cue was part of the row or the column (see a). Here results were similar to the category-learning case, where IWL exhibited catastrophic forgetting when trials were blocked, regardless of rotation condition. IWL also failed to generalize compositionally, failing on the 16 held-out test cues (green lines) in all conditions. (g-i) Experiments using networks capable of both in-weight and in-context learning (ICL). (g) After metalearning, the models again exhibited a blocking advantage, but also showed strong compositional generalization, as shown by the high few-shot test accuracy in the blocked condition. ICL again failed in the rotated condition. (h) After finetuning, the network exhibited an interleaving advantage in the rotated condition, due to catastrophic forgetting when trials were blocked (see i). (i) When trials were blocked in the rule-like condition, accuracy was near-perfect, resulting in little loss and thus little IWL. In the rotated condition, ICL failed, resulting in larger losses, increased IWL, and increased catastrophic forgetting, as can be seen in the rapid drop in accuracy on the first group (‘TrainA,’ shown in purple) while training on the second group (‘TrainB,’ shown in orange), and vice versa. No catastrophic forgetting occurred in the interleaved condition, but compositional generalization (green) was considerably worse than when trials were blocked.
@@
We focused our investigations on a recent study demonstrating compositional generalization in humans on a novel rule-governed task [see Figure 3a-c; 19]. Notably, this study showed that compositional generalization indeed depended on the curriculum, improving when related trials were blocked compared to interleaved—consistent with the idea that the mechanisms underlying compositionality can be linked to those responsible for producing the blocking advantage. This task therefore provides an excellent testbed for our metalearning neural networks, allowing us to replicate our original curriculum-related results from the category-learning task while also studying their connection to compositionality.
@@
In the original task, participants learned to pair colored animals with arbitrary xy-coordinates via trial-and-error. Importantly, the correct locations varied systematically with the two features: color determined the x-coordinate (each of 5 different colors was linked to one of 5 different x-values) while the animal determined the y-coordinate, or vice-versa. Participants saw only 9 of the 25 possible color-animal pairs as study examples; they had to make novel inferences on the 16 remaining pairs during testing (without feedback). This task can be seen as rule-based in that a simple rule (e.g., color = x, animal = y) governs the locations, and can be seen as compositional in that good test performance requires composition of knowledge about a particular color (e.g., ‘blue’ means x = 3) with knowledge about a particular animal (e.g., ‘alligator’ means y = 2) into a novel combination (e.g., ‘blue alligator’ means location is 3, 2).
@@
The key experimental variable manipulated in the study was the curriculum—which 9 of the 25 cues were used as study examples, and the order in which they were presented (see Figure 3a-c). In the Blocked condition, all cues of a particular color (i.e., a single row/column) were presented before all the cues with a particular animal, or vice-versa. In the Interleaved condition, a single row and column were again chosen for study, but their order was randomly shuffled. (Note that the original study also tested two other related conditions, where sampling of items was “Aligned” or “Misaligned”; we simulated these cases and reproduced similar results in the Appendix, but here focus on the key blocked vs. interleaved contrast).
@@
2.2.1 IWL is non-compositional and produces an interleaving advantage
@@
2.2.2 ICL can be compositional and can produce a blocking advantage
We then endowed the network with ICL abilities by metalearning on a distribution of tasks (see methods for details). After metalearning, the network’s emergent ICL algorithm generalized compositionally on unseen tasks, achieving good performance on color-animal combinations that were not included in the study examples. This generalization performance involved the composition of rules that could be inferred from the study examples (see Figure 3b). Furthermore, as in the previous simulations, metalearning endowed the network with an ICL algorithm that exhibits the same kind of blocking advantage observed in humans [19], performing better in the few-shot setting when trials were blocked compared to interleaved (main effect of curriculum on the rule-like task: $p<10^{-3}$).
These findings extend recent work [53] by showing that the ICL algorithm that emerges in metalearning neural networks can reproduce human-like compositional generalization performance and its associated blocking advantage in this experimental paradigm [19]. This is significant because it shows how neural networks, which have traditionally been criticized for lacking compositionality [20, 40], can through metalearning come to implement an ICL algorithm that is capable of human-like compositional generalization [26, 67].
2.2.3 ICL and IWL produce compositionality and both curriculum effects
@@
2.3 LLMs exhibit compositionality and a blocking advantage
@@
Although the datasets used for LLM pretraining are developmentally unrealistic in many ways [70, 71, 72], they are more naturalistic in the sense that they are comprised of natural language text, rather than content that is specifically relevant to our tasks. These corpora are not purposefully designed to encourage ICL or any of our hypothesized properties to emerge. Nevertheless, impressive ICL abilities do arise in these models, giving them the flexibility to accomplish many kinds of tasks in context [48, 49]. Given the scale and complexity of their pretraining datasets, it is unclear a priori what ICL properties LLMs’ should develop, but prior work has shown that their emergent ICL abilities can exhibit compositional generalization in some settings [54, 52, 55], and can also be sensitive to the order in which in-context examples are provided [73, 74].
We thus hypothesized that the properties of ICL assumed by our theoretical framework (i.e., compositionality and a blocking advantage, see principle 2) may emerge in LLMs. We tested this hypothesis with two pretrained LLMs on the same compositional task used above: Llama 2 [75] and GPT-3.5 [48, 76]. We evaluated the emergent ICL abilities of these models by presenting color-animal pairs from the compositional task only in context. [^1] In principle, these models should also show IWL properties like any other neural network, but it is highly expensive to finetune them, and our main questions here pertain to ICL.
@@
Both LLMs showed impressive compositional generalization performance on the task (see Figure 4), even though they were only given the 9 study examples and had not been explicitly pretrained on variants of the task. This shows that the emergent ICL abilities in these models can produce the kinds of generalization behaviors that standard IWL in neural networks struggles to achieve (see test accuracy in Figure 3f).
Notably, both LLMs also produced the blocking advantage in the rule-like version of the task (curriculum main effect: $p<10^{-3}$). [^2] Like our metalearned neural networks, the LLMs also showed the full pattern of curriculum effects described in the human study, see Appendix for details. This again shows that even though the ICL capability in the LLMs has not been specifically sculpted to produce this blocking advantage, it emerges nonetheless via large-scale next-token prediction on large corpora of text.
@@
Thus, neural networks can come to implement an ICL algorithm with the properties of compositionality, a blocking advantage, and a preference for rule-like structure—even when their training does not specifically target these properties, but consists in next-token prediction on a large corpus of natural text.
3 Discussion
@@
Influential theories in cognitive science posit two distinct systems to account for findings suggesting a duality in human learning [1, 2, 3, 4, 5, 6, 7, 8, 9]. Prominent theories leverage distinctions between controlled vs. automatic processing [77, 78, 6], model-based vs. model-free reinforcement learning [3, 79, 80], working memory in PFC vs. striatal synaptic learning [45, 29, 9, 63], system 2 vs. system 1 thinking [33], and rule-based vs. procedural learning [1, 5]. These theories explain why human learning exhibits different properties under different conditions. Here, we have focused on two such properties: 1) compositionality and 2) curriculum effects. Humans are capable of utilizing rule-like structure to generalize compositionally [19, 24, 12, 21, 22, 23], and of integrating over multiple dimensions and making arbitrary associations when no rule-like structure is present [81, 1, 5, 38]. In the former case, learning tends to benefit when related trials are blocked over time [19, 15, 5], while in the latter case it benefits when trials are interleaved [31, 32, 5, 30].
@@
Our work shows how these phenomena can be explained by a single neural network capable of two qualitatively distinct learning processes. In particular, we have shown how metalearning can endow a network with a capacity to learn in context, and how this capacity can capture compositionality and the blocking advantage on tasks governed by rule-like structure. ICL operates by default, but can be unsuccessful on tasks lacking such structure, triggering error-driven IWL and producing an interleaving advantage due to catastrophic forgetting [37, 38]. This dynamic interaction between ICL and IWL is analogous to a tradeoff observed in humans: when working memory is used to learn new stimulus-response rules, reductions in neural prediction errors and incremental reinforcement learning are observed [63, 9]. Our theoretical framework offers a unified perspective on compositionality and curriculum effects, extending dual-process theories by showing how two distinct learning processes can coexist (and compete) within a single neural network.
3.1 Curriculum Effects
@@
There has been some debate about whether humans learn better when related content is blocked or interleaved over time, with some studies finding a blocking advantage [28, 19, 15, 5] and others finding an interleaving advantage [31, 32, 5, 30]. There may be multiple factors that distinguish these cases [e.g., between-category and within-category similarity; 82], but one important variable may be the presence of rule-like structure: humans have been shown to exhibit a blocking advantage when the task is governed by succinct rules, and an interleaving advantage when the task does not afford such rules [19, 5]. These effects are explained by a dual-process account in which a rule-based learning system operates by an explicit hypothesis-testing strategy and a procedural learning system operates by incrementally integrating information over time [1, 5]. Our work offers a novel perspective on this dual-process account, showing how a similar duality can emerge in neural networks capable of both ICL and IWL.
@@
The blocking advantage, on the other hand, doesn’t emerge by default in standard neural networks, but a number of studies have explored the neural mechanisms that might underlie it. For example, a neural network model of rule-based inference and working memory in the PFC showed that blocking related trials over time can encourage abstract rule-like representations to emerge in the network’s activations [14]. More recent work [36] showed that a PFC-like neural network augmented with a gating mechanism and a bias for active maintenance produces a blocking advantage on a task involving cognitive maps [84]. Related work has shown how a neural network equipped with a specialized Hebbian gating mechanism [85] can reproduce a blocking advantage observed in humans on an analogous task [15]. A similar Hebbian mechanism was then used to explain the blocking advantage observed in the compositional task studied here [19]. Another recent study showed how the blocking advantage observed in humans on a next-state prediction task [28] was reproduced by a neural network model that actively maintained distinct contextual representations over time [86]. Overall, these studies emphasize how a blocking advantage can emerge when inferences are made through forward activation dynamics (i.e., in context), such as those made over items maintained in working memory in PFC.
@@
Our theoretical account of the blocking advantage is broadly consistent with previous models of this effect, but has a number of advantages. First, we have shown how the blocking advantage can coexist with the interleaving advantage in a neural network. Furthermore, while our framework is consistent with previous models in suggesting that the blocking advantage is related to activation dynamics [e.g., working memory in PFC; 14, 36], we show how these dynamics can be metalearned in a transformer by training it on a distribution of related tasks [47], thus providing a conceptual link between these prior models and ongoing work investigating metalearning and cognitive flexibility in natural and artificial intelligence [64, 53, 87, 88, 52].
@@
Indeed, we also observed a blocking advantage in LLMs, which have revolutionized artificial intelligence research [48, 49], and are arguably the most cognitively flexible systems ever built [49, 89]. These results show that a blocking advantage can emerge with ICL even when networks are trained on natural text rather than metalearning datasets specifically designed to promote it. Although it is difficult to know exactly why this blocking advantage emerges in the LLMs, we speculate that it is driven by distributional properties of the natural text corpora on which they are trained, such as the tendency for human writing to afford inferences best made by assimilating consecutive examples in a sequential, rather than haphazard, manner. However, further work is needed to better understand the sources of the blocking advantage in the LLMs, and the internal mechanisms responsible for producing it.
@@
In particular, our work does not directly address whether the blocking advantage observed in humans emerges due to strong constraints imposed by neural architecture (e.g., recurrence, limitations in working memory capacity), rather than the statistical properties of the environment (e.g., the distributional properties of natural language). In our experiments, both the metalearning networks and the LLMs utilized the transformer architecture [90], which is not recurrent and does not have hard constraints in working memory capacity. The blocking advantage emerged in these models due to statistical properties of their training data. This was especially clear in the metalearning experiments, where we had full control over the data distribution and confirmed that it determines when the blocking advantage emerges (see Appendix). Consistent with these findings, it has been shown that the human blocking advantage depends on the extent to which the feature dimensions relevant to the rule-like structure of the task are represented in a strongly segregated manner [15], a factor that is likely to depend on an individual’s prior learning experiences. However, we think that the human blocking advantage is also likely to depend on key architectural features of the human brain, such as its recurrence and the mechanisms for gating and serial attention in PFC and basal ganglia [13, 29, 36]. These, in turn, might affect the distributional properties of natural language that is produced by humans and provided as training data for the LLMs. Further work is required to understand how architectural features interact with the distributional properties of a network’s training data, and how they might impact the emergence of ICL with specific properties.
3.2 Compositionality

### v4

@@ Results
Figure 3: Compositional task and results. The task is derived from a human study [20]. Networks were presented with the locations corresponding to particular cues (colored animals) and had to predict the locations of unseen cues. Cues were comprised of sequences of tokens (e.g., ‘blue alligator’). See Appendix for details. (a) Curriculum conditions. In both the blocked and interleaved conditions, the 9 study examples always included one full row and one full column. In the blocked condition, the row was presented in full before the column, or vice-versa. In the interleaved condition, these 9 examples were randomly shuffled. (b) In the rule-like condition, locations varied systematically with the color and animal features (e.g., color determined x-coordinate and animal determined y-coordinate). (c) In our novel rotated condition, the original axes were rotated by 45 degrees, so that any change in either color or animal resulted in a change to both x- and y-coordinates. (d-f) Performance on the compositional task with in-weight learning (IWL) only, where again randomly initialized networks were trained from scratch. (d) The few-shot evaluation tested networks’ ability to solve the task in context based on the 9 study examples given in the input. Again, without prior metalearning neural networks were incapable of solving the task in this way, regardless of condition. (e) Without prior metalearning experience, the network was still able to learn via IWL, performing well on both the rule-like and the rotated tasks after task-specific training. IWL again exhibited an interleaving advantage due to catastrophic forgetting (see f). (f) Accuracy and loss results over the course of IWL training in each of the four conditions. Accuracy is again split by group, in this case corresponding to whether the cue was part of the row or the column (see a). Here results were similar to the category-learning case, where IWL exhibited catastrophic forgetting when trials were blocked, regardless of rotation condition. IWL also failed to generalize compositionally, failing on the 16 held-out test cues (green lines) in all conditions. (g-i) Experiments using networks capable of both in-weight and in-context learning (ICL). (g) After metalearning, the models again exhibited a blocking advantage, but also showed strong compositional generalization, as shown by the high few-shot test accuracy in the blocked condition. ICL again failed in the rotated condition. (h) After task-specific training, the network exhibited an interleaving advantage in the rotated condition, due to catastrophic forgetting when trials were blocked (see i). (i) When trials were blocked in the rule-like condition, accuracy was near-perfect, resulting in little loss and thus little IWL. In the rotated condition, ICL failed, resulting in larger losses, increased IWL, and increased catastrophic forgetting, as can be seen in the rapid drop in accuracy on the first group (‘TrainA,’ shown in purple) while training on the second group (‘TrainB,’ shown in orange), and vice versa. No catastrophic forgetting occurred in the interleaved condition, but compositional generalization (green) was considerably worse than when trials were blocked.
@@
We focused our investigations on a recent study demonstrating compositional generalization in humans on a novel rule-governed task, where the goal was not to categorize stimuli but to learn a latent compositional coordinate system [see Figure 3a-c; 20]. Notably, this study showed that compositional generalization indeed depended on the curriculum, improving when related trials were blocked (47/58 participants generalized) compared to interleaved (36/60 participants generalized)—consistent with the idea that the mechanisms underlying compositionality can be linked to those responsible for producing the blocking advantage. This task therefore provides an excellent testbed for our metalearning neural networks, allowing us to replicate our original curriculum-related results in a different paradigm while also studying their connection to compositionality.
@@
In the original task, participants learned to pair colored animals with arbitrary xy-coordinates via trial-and-error. Importantly, the correct locations varied systematically with the two features: color determined the x-coordinate (each of 5 different colors was linked to one of 5 different x-values) while the animal determined the y-coordinate, or vice-versa. Participants saw only 9 of the 25 possible color-animal pairs as study examples; they had to make novel inferences to generalize to the 16 remaining pairs during testing (without feedback). This task can be seen as rule-based in that a simple rule (e.g., color = x, animal = y) governs the locations, and can be seen as compositional in that good test performance requires composition of knowledge about a particular color (e.g., ‘blue’ means x = 3) with knowledge about a particular animal (e.g., ‘alligator’ means y = 2) into a novel combination (e.g., ‘blue alligator’ means location is 3, 2).
@@
The key experimental variable manipulated in the study was the curriculum—which 9 of the 25 cues were used as study examples, and the order in which they were presented (see Figure 3a). In the Blocked condition, all cues of a particular color (i.e., a single row/column) were presented before all the cues with a particular animal, or vice-versa. In the Interleaved condition, a single row and column were again chosen for study, but their order was randomly shuffled. [^3] Note that the original study also tested two other related conditions, where sampling of items was “Aligned” or “Misaligned”; we simulated these cases and reproduced similar results in the Appendix, but here focus on the key blocked vs. interleaved contrast.
@@
IWL is non-compositional and produces an interleaving advantage
@@
ICL can be compositional and can produce a blocking advantage
We then endowed the network with ICL abilities by having it metalearn on a distribution of tasks (see methods for details). After metalearning, these ICL abilities allowed the network to generalize compositionally on unseen tasks, achieving good performance on color-animal combinations that were not included in the study examples. This generalization performance involved the composition of rules that could be inferred from the study examples (see Figure 3b). Furthermore, as in the previous simulations, ICL exhibited the same kind of blocking advantage observed in humans [20], performing better in the few-shot setting when trials were blocked compared to interleaved (main effect of curriculum on rule-like task: $p<10^{-3}$).
These findings extend recent work [55] by showing that the ICL algorithm that emerges in metalearning neural networks can reproduce human-like compositional generalization behavior and its associated blocking advantage in this experimental paradigm [20]. This is significant because it shows how neural networks, which have traditionally been criticized for lacking compositionality [21, 38], can through metalearning come to implement an ICL algorithm that is capable of human-like compositional generalization [27, 70].
ICL and IWL produce compositionality and both curriculum effects
@@
LLMs exhibit compositionality and a blocking advantage
@@
Although the datasets used for training LLMs are developmentally unrealistic in many ways [73, 74, 75], they are more naturalistic in the sense that they are comprised of natural language text, rather than content that is specifically relevant to our tasks. These corpora are not purposefully designed to encourage ICL or any of our hypothesized properties to emerge. Nevertheless, impressive ICL abilities do arise in these models, giving them the flexibility to accomplish many kinds of tasks in context [50, 51]. Given the scale and complexity of their training datasets, it is unclear a priori what ICL properties LLMs’ should develop, but prior work has shown that their emergent ICL abilities can exhibit compositional generalization in some settings [56, 54, 57], and can also be sensitive to the order in which in-context examples are provided [76, 77].
We thus hypothesized that the properties of ICL assumed by our theoretical framework (i.e., compositionality and a blocking advantage, see principle 2) may emerge in LLMs. We tested this hypothesis by evaluating two LLMs, Llama 2 [78] and GPT-3.5 [50, 79], on the same compositional task used above. We evaluated the emergent ICL abilities of these models by presenting color-animal pairs from the compositional task only in context. [^4] In principle, these models should also show IWL properties like any other neural network, but it is highly expensive to finetune them, and our main questions here pertain to ICL.
@@
Both LLMs showed strong compositional generalization performance on the task (see Figure 4), even though they were only given the 9 study examples and had not been explicitly trained on variants of the task. This shows that the emergent ICL abilities in these models can produce the kinds of generalization behaviors that standard IWL in neural networks struggles to achieve (see test accuracy in Figure 3f).
Notably, both LLMs also produced the blocking advantage in the rule-like version of the task (curriculum main effect: $p<10^{-3}$). [^5] Like our metalearned neural networks, the LLMs also showed the full pattern of curriculum effects described in the human study, see Appendix for details. This again shows that even though the ICL capability in the LLMs has not been specifically sculpted to produce this blocking advantage, it emerges nonetheless via large-scale next-word (next-token) prediction on large corpora of text.
@@
Thus, neural networks can come to implement an ICL algorithm with the properties of compositionality, a blocking advantage, and a preference for rule-like structure—even when their training does not specifically target these properties, but consists in next-token prediction on naturalistic text.
Tradeoff between flexibility and retention
Figure 5: Tradeoff between flexibility and retention. Ablating attention to examples in the context (analogous to limiting the number of items accessible in WM) hurts cognitive flexibility but improves retention. When attention is ablated (i.e., $p_{a}$ is high), few-shot generalization via ICL suffers (left) and incremental learning via IWL is slower (middle left). This results in more errors (middle right), consistent with human EEG data showing prolonged presence of neural signatures of prediction errors under higher WM load [9]. When more errors are made, more IWL occurs, resulting in better retention in the absence of contextual information (right), consistent with human results showing better retention after learning under higher WM load [17, 65]. Results are from the compositional task. $p_{a}$ is the probability that attention to each example in the context was inhibited.
So far, we have highlighted the advantages of ICL over IWL in supporting rapid few-shot learning and generalization, and only noted that IWL is necessary in cases where ICL is less effective (e.g., when trials are interleaved or in rotated tasks). However, an additional benefit of IWL (and relatedly, episodic and semantic memory compared to WM) is that retrieval of information stored in synaptic weights does not require persistent activity (e.g., throughout a delay), and can therefore operate over longer timescales and in the absence of contextual cues [80, 63]. Thus, we hypothesized that a natural consequence of the coexistence of ICL and IWL would be the emergence of a tradeoff between flexibility, or rapid adaptation to new tasks from only a few examples, and retention, or the ability to recall information after longer delays or in the absence of contextual information. The key insight is that if learning takes place in a setting where ICL is successful, generalization based on latent rules may be improved, but fewer prediction errors will update weights. When latent rules are identified, they should give rise to reduced prediction errors, and this suppression should in turn relate to better generalization but worse retention.
In fact, analogous findings have been reported in two lines of research on reinforcement learning (RL) in humans [11, 81, 9, 17, 65]. When the task was structured with a hidden rule that could be inferred from the context, neural reward prediction errors were suppressed, and this suppression predicted better generalization of that rule [11]. In simpler RL tasks, participants also show particularly rapid learning when they can acquire contingencies using an ICL-like WM strategy, achieving near-perfect performance within only a few presentations of each stimulus [9]. However, when WM load was increased, many more presentations were required to achieve the same performance, consistent with an IWL-like incremental learning strategy relying more on RL. EEG recordings showed that neural signatures of reward prediction errors were suppressed when WM load was low [9] or when the underlying structure of the task had been inferred [11]. Moreover, this neural marker of an ICL-like WM strategy was predictive of both better generalization [11] and reduced retention [17]. To date, however, no single model has accounted for both sets of findings.
We tested whether a similar tradeoff between flexibility and retention would emerge in our neural networks. As a proxy for WM load, we ablated the networks’ ability to maintain contextual information throughout learning by inhibiting attention to each in-context example with probability $p_{a}$ (similar results were obtained by adding noise to the attention mechanism itself; see Appendix). Retention was evaluated by testing networks on examples they had seen during training, but in the absence of any contextual information (analogous to testing after a delay when relevant information is not available in WM).
The results for the compositional task are shown in Figure 5 (results for the category-learning task can be found in the Appendix). In both tasks, we observed a robust tradeoff between flexibility and retention. When the ablation was stronger ($p_{a}$ was higher), few-shot generalization via ICL was worse (Figure 5, left) and incremental learning via IWL took longer to reach optimal performance (middle left). This meant that more errors were made throughout learning (middle right), consistent with the stronger neural signatures of prediction errors observed in humans under higher WM load [9, 64]. More errors resulted in greater IWL, leading the model to perform better on the retention test (right), consistent with the superior retention observed in humans that had learned under higher WM load [17, 65].
Thus, the same principles that allow the networks to reproduce compositional generalizations and curriculum effects can also explain the tradeoff observed in human RL experiments [9, 17, 65]. This tradeoff suggests that ICL and IWL have distinct advantages whose relative importance depends on whether flexibility or retention is prioritized.
Discussion
@@
Influential theories in cognitive science posit two distinct systems to account for findings suggesting a duality in human learning [1, 2, 3, 4, 82, 5, 6, 7, 8, 9]. Prominent theories leverage distinctions between controlled vs. automatic processing [83, 84, 6], model-based vs. model-free reinforcement learning [3, 85, 86], WM in PFC vs. striatal synaptic learning [46, 30, 9, 17], system 2 vs. system 1 thinking [34], and rule-based vs. procedural learning [1, 5]. These theories explain why human learning exhibits different phenomena under different conditions. Here, we have focused on three such phenomena: 1) compositionality 2) curriculum effects, and 3) the tradeoff between flexibility and retention. Humans are capable of utilizing rule-like structure to generalize compositionally [20, 25, 12, 22, 23, 24], and of integrating over multiple dimensions and making arbitrary associations when no rule-like structure is present [87, 1, 5, 44]. In the former case, learning can be rapid and flexible, and tends to benefit when related trials are blocked over time [20, 15, 5]. In the latter case, it benefits when trials are interleaved [32, 33, 5, 31], and can result in improved retention but limited flexibility and generalization.
@@
Our work shows how these phenomena can be explained by a single neural network capable of two qualitatively distinct learning processes. In particular, we have shown how metalearning can endow a network with a capacity to learn in context, and how this capacity can capture compositionality and the blocking advantage on tasks governed by rule-like structure. ICL operates when tasks are consistent with readily identifiable rules, but can be unsuccessful on tasks lacking such structure, triggering error-driven IWL and producing an interleaving advantage due to catastrophic forgetting [43, 44]. This dynamic interaction between ICL and IWL naturally recapitulates the tradeoff between flexibility and retention observed in humans: WM can be leveraged to rapidly learn new stimulus-response rules, but causes reductions in neural prediction errors driving incremental reinforcement learning, resulting in worse retention after longer delays [17, 9, 65].
ICL has recently emerged as an important topic of research in machine learning and artificial intelligence [50, 88]. Studies have investigated what kinds of data-distributional properties [58, 63] or neural architectures [89, 90, 91] drive its emergence, as well as the kind of learning algorithm it implements [92, 93, 60], and the internal circuits underlying it [94, 95, 96]. Here, we link this recent work to cognitive science, showing how the dynamic interplay between ICL and IWL can offer a unified perspective on compositionality and curriculum effects. Our work complements prominent dual-process theories by showing how two distinct learning processes can coexist (and compete) within a single neural network.
Curriculum Effects
@@
There has been some debate about whether humans learn better when related content is blocked or interleaved over time, with some studies finding a blocking advantage [29, 20, 15, 5] and others finding an interleaving advantage [32, 33, 5, 31]. There may be multiple factors that distinguish these cases [e.g., between-category and within-category similarity; 97], but one important variable may be the presence of rule-like structure: humans have been shown to exhibit a blocking advantage when the task is governed by succinct rules, and an interleaving advantage when the task does not afford such rules [20, 5]. These effects are explained by a dual-process account in which a rule-based learning system operates by an explicit hypothesis-testing strategy and a procedural learning system operates by incrementally integrating information over time [1, 5]. Our work offers a novel perspective on this dual-process account, showing how a similar duality can emerge in neural networks capable of both ICL and IWL.
@@
The blocking advantage, on the other hand, doesn’t emerge by default in standard neural networks, but a number of studies have explored the neural mechanisms that might underlie it. For example, a neural network model of rule-based inference and WM in the PFC showed that blocking related trials over time can encourage abstract rule-like representations to emerge in the network’s activations [14]. More recent work [42] showed that a PFC-like neural network augmented with a gating mechanism and a bias for active maintenance produces a blocking advantage on a task involving cognitive maps [99]. Related work has shown how a neural network equipped with a specialized Hebbian gating mechanism [100] can reproduce a blocking advantage observed in humans on an analogous task [15]. A similar Hebbian mechanism was then used to explain the blocking advantage observed in the compositional task studied here [20]. Another recent study showed how the blocking advantage observed in humans on a next-state prediction task [29] was reproduced by a neural network model that actively maintained distinct contextual representations over time [101]. Overall, these studies emphasize how a blocking advantage can emerge when inferences are made through forward activation dynamics (i.e., in context), such as those made over items maintained in WM in PFC.
@@
Our theoretical account of the blocking advantage is broadly consistent with previous models of this effect, but has a number of advantages. First, we have shown how it can emerge in a neural network model that also produces the interleaving adantage on different tasks. Furthermore, while our framework is consistent with previous models in suggesting that the blocking advantage is related to activation dynamics [e.g., WM in PFC; 14, 42], we show how these dynamics can be metalearned by training on a distribution of related tasks [49], thus providing a conceptual link between these prior models and ongoing work investigating metalearning and cognitive flexibility in natural and artificial intelligence [66, 55, 68, 102, 103, 54].
@@
Indeed, we also observed a blocking advantage in LLMs, which have revolutionized artificial intelligence research [50, 51] and appear to exhibit high levels of cognitive flexibility [51, 104]. These results show that a blocking advantage can emerge with ICL even when networks are trained on natural text rather than metalearning datasets specifically designed to promote it. Although it is difficult to know exactly why this blocking advantage emerges in the LLMs, we speculate that it is driven by distributional properties of the natural text corpora on which they are trained, such as the tendency for human writing to afford inferences best made by assimilating consecutive examples in a sequential, rather than haphazard, manner. However, further work is needed to better understand the sources of the blocking advantage in the LLMs, and the internal mechanisms responsible for producing it.
@@
In general, our work does not directly address whether the blocking advantage observed in humans emerges due to strong constraints imposed by neural architecture (e.g., recurrence, limitations in WM capacity), rather than the statistical properties of the environment (e.g., the distributional properties of natural language). In our experiments, the metalearning networks and the LLMs utilized the transformer architecture [105], which is not recurrent and does not have hard constraints in WM capacity. Both the blocking advantage and the preference for rule-like tasks emerged in these models due to the statistical properties of their training data. This was especially clear in the metalearning experiments, where we had full control over the data distribution and confirmed that it determines when the blocking advantage emerges (see Appendix). Consistent with these findings, prior work has shown that metalearning networks trained on category-learning problems that match the natural statistics of real-world tasks perform poorly on the same problems that humans struggle with [68]. Furthermore, the human blocking advantage has been shown to depend on the extent to which the feature dimensions relevant to the rule-like structure of the task are represented in a strongly segregated manner [15], a factor that is likely to depend on an individual’s prior learning experiences. However, we think that the human blocking advantage is also likely to depend on key architectural features of the human brain, such as its recurrence and the mechanisms for gating and serial attention in PFC and basal ganglia [13, 30, 14, 42]. These, in turn, might affect the distributional properties of natural language that is produced by humans and provided as training data for the LLMs [106]. Further work is required to understand how architectural features interact with the distributional properties of a network’s training data, and how they might impact the emergence of ICL with specific properties.
Compositionality